In recent years, Novak Transportation purchased three used buses. Because of frequent turnover in the accounting department, a different accountant selected the depreciation method for each bus, and various methods were selected. Information concerning the buses is summarized as follows.

Bus

Acquired

Cost

Salvage
Value

Useful Life
in Years

Depreciation
Method

1 1/1/18 $ 99,000 $ 5,500 5 Straight-line
2 1/1/18 110,000 10,500 4 Declining-balance
3 1/1/19 89,000 8,500 4 Units-of-activity

For the declining-balance method, the company uses the double-declining rate. For the units-of-activity method, total miles are expected to be 115,000. Actual miles of use in the first 3 years were 2019, 23,000; 2020, 31,000; and 2021, 30,500.

For Bus #3, calculate depreciation expense per mile under units-of-activity method. (Round answer to 2 decimal places, e.g. 0.50.)

Depreciation expense $

per mile

then Compute the amount of accumulated depreciation on each bus at December 31, 2019.
Answer #1

Answer- Depreciation expense per mile under units-of-activity method=$0.70 per mile.

Explanation- Activity based depreciation method:-Annual depreciation expense per unit-

=Cost – salvage /Total miles

=($89000-$8500)/115000 miles

=$0.70 per mile

Answer- The amount of accumulated depreciation on Bus 1 at December 31, 2019 = $18700+$18700

= $37400

Explanation- Straight line Method-

= Cost of asset- Salvage value of asset/No. of useful life (years)

=($99000-$5500)/5 years

=$93500/5 years

= $18700

Annual depreciation expense= $18700  

Answer- The amount of accumulated depreciation on Bus 2 at December 31, 2019 =$27500+$55000

= $82500.

Explanation-

Double Declining balance depreciation is calculated using the following formula=

Depreciation = Depreciation Rate * Book Value of Asset

Depreciation rate is given by the following formula:

Depreciation Rate = Accelerator *Straight Line Rate

Straight-line Depreciation Rate = 1/4 = 0.25 = 25%
Declining Balance Rate = 2*25% = 50%

Depreciation for Year 2018 = $110000*50% =55000

Book value at the end of year 2018 = $110000-$55000 =$55000

Depreciation for Year 2019 = $55000*50% = $27500

Answer- The amount of accumulated depreciation on Bus 3 at December 31, 2019 =$16100.

Explanation- Activity based depreciation method:-Annual depreciation expense per unit-

=Cost – salvage /Total miles

=($89000-$8500)/115000 miles

=$0.70 per mile

Depreciation expense in year 2019 = Depreciation expense per mile* Actual miles

=$0.70 per mile*23000 miles

=$16100
Answer #2

BUS -1 Cost of Asset 99000 Residual Value 5500 Useful Life Depreciable Asset = Cost - Residual value Depreciable Asset 93500BUS-2 Depreciation rate = ( 1*200 /useful life ) = 50% Double Decline Method Year 2018 2019 2020 Total Book value at BeginninBUS -3 Cost of Asset 89000 Residual Value 8500 Total production/hours 115000 Depreciable Asset = Cost - Residual value Deprec